The probability of a randomly selected adult in one country being infected with a certain virus was 0.004. In tests for the virus, blood samples from 27 people are combined.

What is the probability that the combined sample tests positive for the virus?

Is it unlikely for such a combined sample to test positive? Note that the combined sample tests positive if at least one person has the virus. The probability that the combined sample will test positive is?
Answer #1

1)

probability that the combined sample tests positive for the virus =1-P(none is infected with virus)

=1-(1-0.004)27=0.102567

it is not unlikely as the probability of this event is higher then 0.05 level